# MERA Code: A Unified Framework for Evaluating Code Generation Across Tasks

### v2

@@ Abstract
Advancements in LLMs have enhanced task automation in software engineering; however, current evaluations primarily focus on natural language tasks, overlooking code quality. Most benchmarks prioritize high-level reasoning over executable code and real-world performance, leaving gaps in understanding true capabilities and risks associated with these models in production. To address this issue, we propose MERA Code, a new addition to the MERA benchmark family, specifically focused on evaluating code for the latest code generation LLMs in Russian. This benchmark includes 11 evaluation tasks that span 8 programming languages. Our proposed evaluation methodology features a taxonomy that outlines the practical coding skills necessary for models to complete these tasks. The benchmark comprises an open-source codebase for users to conduct MERA assessments, a scoring system compatible with various programming environments, and a platform featuring a leaderboard and submission system. We evaluate open LLMs and frontier API models, analyzing their limitations in terms of practical coding tasks in non-English languages. We are publicly releasing MERA to guide future research, anticipate groundbreaking features in model development, and standardize evaluation procedures.
@@
Artem Chervyakov1, Alexander Kharitonov1, Pavel Zadorozhny1, Adamenko Pavel1, Rodion Levichev1, Dmitrii Vorobev1, Dmitrii Salikhov1, Aidar Valeev1,8, Alena Pestova3, Maria Dziuba2,3, Ilseyar Alimova7, Artem Zavgorodnev4, Aleksandr Medvedev4, Stanislav Moiseev4, Elena Bruches6, Daniil Grebenkin6, Roman Derunets6, Vikulov Vladimir5, Anton Emelyanov1, Vladimir V. Ivanov8, Dmitry Babayev1, Valentin Malykh3 Alena Fenogenova1 1 SberAI, 2 ITMO University, 3 MWS AI, 4 T-Technologies, 5 Rostelecom, 6 Siberian Neuronets, 7 Skoltech, 8 Innopolis University Correspondence: mera@a-ai.ru
@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated significant potential for automating software engineering tasks, such as code generation and documentation. However, current evaluation methods predominantly emphasize natural language understanding, often neglecting critical factors like code quality, real-world applicability, and multilingual support. Benchmarks such as HumanEval-X Zheng et al. (2023b), MultiPL-E Cassano et al. (2022), and mxEval Athiwaratkun et al. (2023) address multilingualism through translated datasets but mainly focus on programming languages. They do not fully capture the interaction between natural language (such as requirements and comments) and code, which is vital for real-world development. This limitation restricts our understanding of LLMs’ practical capabilities, particularly in non-English contexts. Tasks such as writing localized documentation or interpreting vague requirements necessitate proficiency in both natural and programming languages, yet no comprehensive benchmark currently evaluates these cross-disciplinary skills. To fill this gap, we introduce MERA Code — a comprehensive evaluation system explicitly designed for the Russian language. MERA Code provides tools to assess LLMs in realistic, practical, multilingual software development scenarios.
@@
  • A reproducible evaluation methodology for LLMs in Russian;
@@
  • An open evaluation platform with a scoring system, framework  [^1] The evaluation code https://github.com/MERA-Evaluation/MERA_CODE released under MIT License, and public leaderboard;
  • A set of performance analyses ranging from open-source general models to proprietary coding APIs.
@@ 2 Related Work
Early benchmarks for evaluating LLM coding capabilities targeted simple, single-function problems with automated test-execution scoring, as exemplified by HumanEval Chen et al. (2021a) and MBPP Austin et al. (2021a), and by datasets drawn from online platforms such as APPS Hendrycks et al. (2021), CodeContests Li et al. (2022a), and their extension in TACO Li et al. (2023). As model context windows expanded, benchmarks introduced tasks of increasing scope: ClassEval Du et al. (2023) required correct implementations at the class level, while CoderEval Du et al. (2023) and RepoBench Liu et al. (2023b) assessed end-to-end project and repository proficiency. To better mirror developer workflows, SWE-Bench Jimenez et al. (2024) curated tasks from real GitHub issues, and BigCodeBench evaluated LLMs’ ability to orchestrate multiple function calls in practical programming scenarios. However, widespread dataset reuse has engendered significant data leakage, prompting the creation of dynamic evaluation suites such as LiveCodeBench Jain et al. (2024) and CodeElo Quan et al. (2025), which, despite mitigating test-set contamination, remain distant from the complexity of industrial software development.
Most existing benchmarks concentrate predominantly on code generation, overlooking the broader spectrum of development tasks such as code repair, execution, and test output prediction that are integral to real-world programming. Initiatives like CodeXGLUE Lu et al. (2021) span fourteen datasets covering ten tasks from defect detection to summarization and translation, and LongCodeArena Bogomolov et al. (2024) further incorporates self-repair and runtime evaluation. While multilingual programming benchmarks (e.g., HumanEval-X Zheng et al. (2023b), MultiPL-E Cassano et al. (2022), mxEval Athiwaratkun et al. (2023)) address various codebases, they generally neglect the multilingual natural-language elements of software engineering—requirements elicitation, documentation, commenting, and review. Consequently, no existing benchmark comprehensively measures LLM performance across the full software development lifecycle and in the diverse natural and programming languages encountered in practice.
@@ 3.1 Overview
MERA Code introduces a practical evaluation framework that expands the Family of MERA benchmarks [^2] https://mera.a-ai.ru/en/code Fenogenova et al. (2024) for assessing code-oriented and general-purpose state-of-the-art (SOTA) models. This framework serves as a foundational resource for both the research and industrial communities, fostering collaboration to improve task coverage and adapt to the evolving capabilities of LLMs. The release includes the proposed taxonomy of the model’s skills needed to solve practical coding tasks, along with a set of public and private test tasks in instruction-based format, an open evaluation codebase, and a web platform featuring an open automatic scoring system and leaderboard.
@@ 3.2.1 Taxonomy
Several works proposed various taxonomies to systematize the field of software development and engineering. In Zhang et al. (2023), the taxonomy aligns with the software development pipeline and considers tasks that must be solved to deliver the product. The work Ruf et al. (2015) analyzes the field of software engineering from a perspective of programming skills required to solve specific tasks.
Our taxonomy is based on the latter approach since it allows us to decompose arbitrary tasks into a limited number of skills. Hence, the resulting taxonomy could be exhaustive yet straightforward.
A language model can be considered as a system consisting of inputs, internal state, and outputs. According to these parts, we derive four fundamental skills as a ground of our taxonomy: Perception (input), Reasoning and Knowledge (internal state), and Generation (output). All other skills form a tree, becoming increasingly niche with each new level. The taxonomy  [^3] We recognize that the current tasks in MERA do not encompass the complete taxonomy of skills. This serves as a guide for the community to understand what is covered and what needs to be added to future tasks in a systematic manner. presented in Figure 1 is based on Ruf et al. (2015), Zhang et al. (2023), and the authors’ domain expertise.
Figure 1: Taxonomy of MERA Code encompassing four foundational skills Perception / Knowledge / Reasoning / Generation utilized in the model to address certain tasks. Detailed explanation of each skill could be found in Appendix B.
@@ 3.2.2 Prompts selection
LLMs’ performance might significantly depend on a given prompt Shin et al. (2020), Gao et al. (2021), Fenogenova et al. (2024). To ensure an unbiased and robust assessment, each of the 11 tasks is paired with a set of distinct prompts. These prompts differ in syntactic framing, level of detail in the task description, and specification of the required output format. Prompts are assigned uniformly at random across all samples for a given task, with exactly one prompt applied per sample and the assignment is fixed. This strategy mitigates any inadvertent advantage or disadvantage conferred by a particular prompt style or model family affinity. Every prompt comprises a clear template for the expected output, thereby aligning model responses with the automated scoring system. The examples of prompts for each task could be found in Appendix A.
@@ 3.2.3 Evaluation procedure
All MERA Code tasks employ purely generative and instructive assessment where models continuously emit tokens until meeting predefined stopping conditions tuned to balance creativity, determinism, and safety. Each raw output undergoes task-specific post-processing to conform with metric requirements, including mandatory extraction of markdown-fenced code blocks from responses (with fallback to full outputs when parsing fails) before being scored with metrics. MERA Code utilizes the following metrics:
Pass@k Chen et al. (2021a) evaluates the functional correctness of the generated code.
Compile@k - evaluates the correctness of the compilation of the generated code.
chrF Popović (2015) is used in code-to-text tasks since its enhanced sensitivity to Russian morphological complexity.
BLEU Papineni et al. (2002a) measures the n-grams level similarity of the prediction and gold answer.
CodeBLEU (CBLEU) Ren et al. (2020a) combines regular BLEU with the measure of similarity of code syntax via abstract syntax trees (AST) and code semantics via data-flow.
Judge@k measures whether one of the top-k predictions matches with the gold answer via LLM-as-a-Judge Zheng et al. (2023a).
@@ 3.3 Tasks
The MERA Code currently encompasses 11 code tasks. The statistics are presented in Table 1. We will briefly discuss the tasks listed below. Additional details and examples can be found in Appendix A.
@@
is a benchmark for evaluating code generation models on the task of producing executable JUnit 5 unit tests for real-world Java classes. The benchmark consists of 227 tasks mined from open-source GitHub repositories. Each task provides the source code of a focal Java class, its name, and the expected test class name; the model must generate a complete and executable test file. Evaluation is fully automated: compile@1 measures the proportion of generations that compile successfully, while pass@1 quantifies those that pass all tests when executed with mvn test inside a provided Docker environment. We also use repotestfootnotemark: to run the tests and measure correctness.
@@
Dziuba and Malykh (2025) is the first benchmark for evaluating LLMs’ ability to generate structured Russian-language code documentation. It comprises 153K examples across Python, Java, JavaScript, C#, and Go, sourced from human-authored GitHub repositories and synthetically generated examples. The $500$ samples were carefully selected from these examples as a test set for MERA Code. Evaluation employs the chrF metric due to its sensitivity to Russian morphological complexity.
@@
Valeev et al. (2025) is a benchmark that evaluates the performance of LLMs on large repositories. It focuses on C and C++ languages since other benchmarks do not fully cover them. YABLoCo consists of $208$ carefully selected functions from four large open-source repositories (bullet3, Redis, OpenSSL, LLVM). Each function is assigned a context category determined by the function’s dependencies. The primary purpose of the benchmark is to evaluate the quality of LLMs’ generations when various repository context is given alongside the prompt. The quality of the generated code is measured by the pass@1 and EM metrics.
@@
The first benchmark for automated code review comment generation in Russian is designed to address the limitations of English-centric and often noisy datasets. The benchmark is built from a high-quality dataset of 689 merge-request diff-comment pairs sourced from an industry-oriented educational program, where professional developers review student code (Java, Python, Go, Scala). A rigorous two-stage curation process is employed, involving an LLM-assisted filter followed by verification from two independent human experts, to ensure that every comment is self-contained, actionable, and reproducible from the code diff alone. As an evaluation metrics, we use Judge@k, chrF, BLEU in a reference-based setup.
@@
is a multilingual dataset for evaluating model abilities to generate unit tests for functions and methods in five programming languages: Python, Java, Go, JavaScript, and C#. It is built from GitHub repositories with open licenses. Each observation consists of a focal function/method, a unit test for it (test function/method), as well as the focal and test contexts collected from the repository for the focal and test functions, correspondingly. The goal of the model is to generate a unit test based on the provided focal function/method and its context, as well as the test function/method context. The evaluation metric is CBLEU.
@@
Private ruCodeEval Python pass@k 164 10 Instruction Following, Code Perception, Completion, Algorithms & Data Structures
@@
Public ruHumanEval Python pass@k 164 10 Instruction Following, Code Perception, Completion
@@ 3.4 System Demo
The automatic scoring system is accessible on the benchmark platform  [^6] https://mera.a-ai.ru/en/code. The submission process involves the following steps. First, users must clone the MERA Code benchmark repository  [^7] The MERA Code is based on the fork of the LM-evaluation harness Gao et al. (2024). and create submission files using a shell script and the provided customized evaluation code. Second, they must register on the benchmark platform and upload their submission files through the interface in their account for automatic evaluation. The scoring process may take up to three hours. Once the evaluation is complete, the results are displayed in the user’s account and remain private unless the user opts to publish them using the “Publish” function. In this case, the submission undergoes expert verification for reproducibility, which includes checking the log files automatically generated by the evaluation script and reviewing the provided submission information. Once approved, the model’s score is publicly visible on the leaderboard, while the specific outputs remain private. The user process for the benchmark is illustrated in Figure 2 of the Appendix.
4 Evaluation
The MERA Code evaluation procedure is designed to strike a balance between accessibility and rigorous, standardized assessment of code generation capabilities. Users begin by cloning the benchmark’s GitHub repository, built on the lm-evaluation-harness framework  [^8] https://github.com/EleutherAI/lm-evaluation-harness with officially integrated customizations, and execute a dedicated script with fixed hyperparameters to launch lm-eval for a specified model and task set. Two repository branches are available: one enables local scoring of public tasks with reference outputs, while the other excludes complex dependencies to facilitate submission to the automated scoring system. While users may choose either workflow, local scoring does not guarantee metric correctness for tasks dependent on runtime checks or hidden test cases. Generated outputs are saved in JSON files and submitted as a ZIP archive via the MERA Code website, along with additional information to ensure reproducibility. The automated system then processes submissions, performing runtime evaluations that may take hours, and delivers detailed metrics to the user’s account page.
4.1 Baselines
We evaluate both code-oriented and general-purpose open models of various sizes: Qwen2.5 Coder (32B-Instruct), Qwen2.5 (72B-Instruct), Deepseek Coder V2 (236B-Instruct), ByteDance Seed-Coder (8B-Instruct), Mixtral (8x22B-Instruct), Yi Coder (9B-Chat), and Vikhr-YandexGPT-5-Lite (8B-it). Additionally, we assess proprietary models such as OpenAI GPT-4  [^9] https://openai.com/api/, Gemini 2.5  [^10] https://deepmind.google/models/gemini/, GigaChat 2 Max  [^11] https://giga.chat/, GigaCode  [^12] https://gigacode.ru/.
4.2 Results
Model Total Score Private Score

### v3

@@ Abstract
Advancements in LLMs have enhanced task automation in software engineering; however, current evaluations primarily focus on natural language tasks, overlooking code quality. Most benchmarks prioritize high-level reasoning over executable code and real-world performance, leaving gaps in understanding true capabilities and risks associated with these models in production. To address this issue, we propose MERA Code, a new addition to the MERA benchmark family, specifically focused on evaluating code for the latest code generation LLMs. This benchmark includes 11 evaluation tasks that span 8 programming languages. Our proposed evaluation methodology features a taxonomy that outlines the practical coding skills necessary for models to complete these tasks. The benchmark comprises an open-source codebase for users to conduct MERA assessments, a scoring system compatible with various programming environments, and a platform featuring a leaderboard and submission system. We evaluate open LLMs and frontier API models, analyzing their limitations regarding practical coding tasks in non-English languages. We are publicly releasing MERA to guide future research, anticipate groundbreaking features in model development, and standardize evaluation procedures.
@@
Artem Chervyakov1, Alexander Kharitonov1, Pavel Zadorozhny1, Pavel Adamenko1, Rodion Levichev1, Dmitrii Vorobev1, Dmitrii Salikhov1, Aidar Valeev1,8, Alena Pestova3, Maria Dziuba2,3, Ilseyar Alimova7, Artem Zavgorodnev4, Aleksandr Medvedev4, Stanislav Moiseev4, Elena Bruches6, Daniil Grebenkin6, Roman Derunets6, Vikulov Vladimir5, Anton Emelyanov1, Vladimir V. Ivanov8, Dmitry Babayev1, Valentin Malykh3 Alena Fenogenova1 1 SberAI, 2 ITMO University, 3 MWS AI, 4 T-Technologies, 5 Rostelecom, 6 Siberian Neuronets, 7 Skoltech, 8 Innopolis University Correspondence: mera@a-ai.ru
@@ 1 Introduction
Recent advances in large language models (LLMs) have demonstrated significant potential for automating software engineering tasks, such as code generation and documentation. However, current evaluation methods predominantly emphasize natural language understanding, often neglecting critical factors such as code quality, real-world applicability, and multilingual support. Benchmarks such as HumanEval-X zheng2023codegeex, MultiPL-E cassano2022multipl, and mxEval athiwaratkun2023multi address multilingualism through translated datasets but mainly focus on programming languages. They do not fully capture the interaction between natural language (such as requirements and comments) and code, which is vital for real-world development. This limitation restricts our understanding of LLMs’ practical capabilities, particularly in non-English contexts. Tasks such as writing localized documentation or interpreting vague requirements necessitate proficiency in both natural and programming languages, yet no comprehensive benchmark currently evaluates these cross-disciplinary skills. To fill this gap, we introduce MERA Code — a comprehensive evaluation system explicitly designed for the Russian language. MERA Code provides tools to assess LLMs in realistic, practical, multilingual software development scenarios.
@@
  • A reproducible evaluation methodology for LLMs;
@@
  • An open evaluation platform with a scoring system, framework  [^1] The evaluation code https://github.com/MERA-Evaluation/MERA_CODE is released under the MIT License, and public leaderboard;
  • A comprehensive evaluation of models performance, ranging from open-source general models to proprietary coding APIs.
@@ 2 Related Work
Early benchmarks for evaluating LLM coding skills focused on short problems scored by automated test‐execution systems, such as HumanEval chen2021evaluating and MBPP austin2021program, and tasks from online platforms like APPS hendrycksapps2021 and CodeContests li2022competition. As context windows grew, evaluations expanded from class-level (ClassEval du2023classeval) to repository-scale tasks (RepoBench liurepobench, SWE-Bench jimenez2024swe), with dynamic suites (LiveCodeBench jainlivecodebench, CodeElo quan2025codeelo) emerging to mitigate data leakage, though they still lack industrial complexity. More fundamentally, the field predominantly privileges code generation, overlooking critical tasks such as code repair, execution, and test output prediction that are partially addressed by broader initiatives (CodeXGLUE lu2021codexglue, LongCodeArena bogomolov2024long). Furthermore, while multilingual programming benchmarks (MultiPL-E cassano2022multipl, HumanEval-X zheng2023codegeex) expand code language coverage, they omit the essential multilingual natural-language aspects of software engineering—such as requirements, documentation, and review. Consequently, no existing benchmark comprehensively evaluates LLMs across the full software development lifecycle in both natural and programming languages as encountered in practice.
@@ 3.1 Overview
MERA Code introduces a practical evaluation framework that expands the Family of MERA benchmarks [^2] https://mera.a-ai.ru/en/code fenogenova-etal-2024-mera for assessing code-oriented and general-purpose state-of-the-art (SOTA) models. This framework serves as a foundational resource for both the research and industrial communities, fostering collaboration to improve task coverage and adapt to the evolving capabilities of LLMs. The release includes the proposed taxonomy of the model’s skills needed to solve practical coding tasks, along with a set of public and private test tasks in an instruction-based format, an open evaluation codebase, and a web platform featuring an open automatic scoring system and leaderboard.
@@ 3.2.1 Taxonomy
Several works proposed various taxonomies to systematize the field of software development and engineering. In zhang2023unifying, the taxonomy aligns with the software development pipeline and considers tasks that must be solved to deliver the product. The work ruf2015classification analyzes the field of software engineering from a perspective of programming skills required to solve specific tasks. Our taxonomy is based on the latter approach since it allows us to decompose arbitrary tasks into a limited number of skills. Hence, the resulting taxonomy could be exhaustive yet straightforward.
A language model can be considered as a system consisting of inputs, internal state, and outputs. According to these parts, we derive four fundamental skills as a ground of our taxonomy: Perception (input), Reasoning and Knowledge (internal state), and Generation (output). All other skills form a tree, becoming increasingly niche with each new level. The taxonomy  [^3] We recognize that the current tasks in MERA do not encompass the complete taxonomy of skills. This serves as a guide for the community to understand what is covered and what needs to be added to future tasks in a systematic manner. presented in Figure 1 is based on ruf2015classification, zhang2023unifying, and the authors’ domain expertise.
Figure 1: Taxonomy of MERA Code encompassing four foundational skills Perception / Knowledge / Reasoning / Generation utilized in the model to address certain tasks. Detailed explanation of each skill could be found in Appendix D.
@@ 3.2.2 Prompts selection
LLMs’ performance might significantly depend on a given prompt shin-etal-2020-autoprompt, gao-etal-2021-making, fenogenova-etal-2024-mera. To ensure an unbiased and robust assessment, each of the 11 tasks is paired with a set of distinct prompts. These prompts differ in syntactic framing, level of detail in the task description, and specification of the required output format. Prompts are assigned uniformly at random across all samples for a given task, with exactly one prompt applied per sample and the assignment is fixed. This strategy mitigates any inadvertent advantage or disadvantage conferred by a particular prompt style or model family affinity. Every prompt comprises a clear template for the expected output, thereby aligning model responses with the automated scoring system. Example prompts for each task are provided in Appendix C. The benchmark is designed to evaluate performance on tasks relevant to Russian-speaking users, and as such, the prompt examples reflect real-world use cases. An ablation study B provides evidence of the language-robust benchmark methodology based on the comparison of the Russian and English instructions.
3.2.3 The impact of the prompts language
Table 1 reports the score differences between Russian- and English-prompted runs [^4] The RuCodeReviewer, ruCodeEval, ruHumanEval, and StRuCom are excluded from the comparison, as the tasks are designed to have natural language as a part of the input data or targets (not only the instruction formulations as in the other benchmark tasks). as median [Q1, Q3] across tasks (scores are bounded between 0 and 1). Overall, the medians are very close to zero, and the interquartile ranges are typically well below 0.1 in absolute magnitude, indicating that prompt language has little practical effect on performance for the benchmark. For most tasks, the IQR straddles or stays near zero (e.g., CodeCorrectness, UnitTests, RealCode), suggesting no consistent advantage of either language, with only a small positive shift on YABLoCo and CodeLinterEval, and a mild negative tendency on RealCodeJava and JavaTestGen. Taken together, these results support the view that the benchmark is largely language-robust (at least between English and Russian); since the tasks are code-centric and much of the input is programming language rather than natural language, models appear to rely primarily on the code and formal constraints, making the evaluation methodology relatively insensitive to the human-language surface form of the instruction (see the detailed comparison in B).
Task Difference
CodeCorrectness (EM) 0.01 [-0.02, 0.04]
CodeLinterEval (pass@1) 0.03 [-0.01, 0.05]
JavaTestGen (pass@1) -0.01 [-0.02, 0.01]
RealCode (pass@1) 0.0 [-0.01, 0.04]
RealCodeJava (pass@1) -0.02 [-0.05, 0.01]
UnitTests (CodeBLEU) 0.0 [-0.01, 0.0]
YABLoCo (pass@1) 0.06 [0.01, 0.07]
Table 1: The statistics on the difference between the scores of the models with the Russian prompts and the English ones. The difference is displayed as "Median [Q1, Q3]". Negative values mean that the score with the English prompts is larger than that with the Russian prompts.
@@ 3.2.4 Evaluation procedure
All MERA Code tasks employ purely generative and instructive assessment where models continuously emit tokens until meeting predefined stopping conditions. Each raw output undergoes task-specific post-processing to conform with metric requirements, including mandatory extraction of markdown-fenced code blocks from responses (with fallback to full outputs when parsing fails) before being scored with metrics. All metrics utilized in MERA Code are listed in Table 2 and described in AppendixA.
@@ 3.3 Tasks
The MERA Code currently encompasses 11 code tasks. The statistics are presented in Table 2. We will briefly discuss the tasks listed below. Additional details and examples can be found in Appendix C.
@@
is a benchmark for evaluating code generation models on the task of producing executable JUnit 5 unit tests for real-world Java classes. The benchmark consists of 227 tasks mined from open-source GitHub repositories. Each task provides the source code of a focal Java class, its name, and the expected test class name; the model must generate a complete and executable test file. Evaluation is fully automated: compile@1 measures the proportion of generations that compile successfully, while pass@1 quantifies those that pass all tests when executed with mvn test inside a provided Docker environment. We also release and use repotestfootnotemark: to run the tests and measure correctness.
@@
dziuba2025strucom is the first benchmark for evaluating LLMs’ ability to generate structured Russian-language code documentation. It comprises 153K examples across Python, Java, JavaScript, C#, and Go, sourced from human-authored GitHub repositories and synthetically generated examples. The $500$ samples were carefully selected from these examples as a test set for MERA Code. Evaluation employs the chrF metric due to its sensitivity to Russian morphological complexity.
@@
valeev2025yabloco is a benchmark that evaluates the performance of LLMs on large repositories. It focuses on C and C++ languages since other benchmarks do not fully cover them. YABLoCo consists of $208$ carefully selected functions from four large open-source repositories (bullet3, Redis, OpenSSL, LLVM). Each function is assigned a context category determined by the function’s dependencies. The primary purpose of the benchmark is to evaluate the quality of LLMs’ generations when various repository contexts are given alongside the prompt. The quality of the generated code is measured by the pass@1 and EM metrics.
@@
The first benchmark for automated code review comment generation in Russian is designed to address the limitations of English-centric and often noisy datasets. The benchmark is built from a high-quality dataset of 689 merge-request diff-comment pairs sourced from an industry-oriented educational program, where professional developers review student code (Java, Python, Go, Scala). A two-stage curation process is employed, involving an LLM-assisted filter followed by verification from two independent human experts, to ensure that every comment is self-contained, actionable, and reproducible from the code diff alone. As evaluation metrics, we use Judge@k, chrF, BLEU in a reference-based setup. The LLM judge is validated with the help of human experts and achieves a Spearman correlation of 0.831 with human annotators, confirming its reliability for the assessment of semantic equivalence. Details about LLM-as-a-Judge validation could be found in Appendix C.
@@
is a multilingual dataset for evaluating model abilities to generate unit tests for functions and methods in five programming languages: Python, Java, Go, JavaScript, and C#. It is built from GitHub repositories with open licenses. Each observation consists of a focal function/method, a unit test for it (test function/method), as well as the focal and test contexts collected from the repository for the focal and test functions, respectively. The goal of the model is to generate a unit test based on the provided focal function/method and its context, as well as the test function/method context. The evaluation metric is CBLEU.
@@
Private
ruCodeEval Python pass@k 164 10 Instruction Following, Code Perception, Completion, Algorithms & Data Structures
@@
Public
ruHumanEval Python pass@k 164 10 Instruction Following, Code Perception, Completion
@@ 3.4 System Demo
The automatic scoring system is accessible on the benchmark platform  [^7] https://mera.a-ai.ru/en/code. The submission process involves the following steps.
First, users must clone the MERA Code benchmark repository  [^8] The MERA Code is based on the fork of the LM-evaluation harness eval-harness. and create submission files using a shell script and the provided evaluation code. Two repository branches are available: one enables local scoring of public tasks with reference outputs, while the other excludes complex dependencies to facilitate submission to the automated scoring system.
Second, they must register on the benchmark platform and upload their submission files through the interface in their account for automatic evaluation. The scoring process may take up to three hours. Once the evaluation is complete, the results are displayed in the user’s account and remain private unless the user opts to publish them. In this case, the submission undergoes expert verification, which includes checking the log files automatically generated by the evaluation script and reviewing the provided submission information. Once approved, the model’s score is publicly visible on the leaderboard, while the specific outputs remain private. The user process for the benchmark is illustrated in Figure 2 of the Appendix.
4 Evaluation results and analysis
Model Citation
Params,
B
Total
Score
Private
Score